OpenStudy (anonymous):

Multiply. (square root of 6 -15th root of 3)(2nd root of 6 + 6th root of 3) Simplify answer. Type exact answer using radicals as needed.

OpenStudy (anonymous):

just foil like any other problem

OpenStudy (anonymous):

\[(\sqrt{6}-\sqrt[15]{3})(\sqrt[2]{6}+\sqrt[6]{3})?\]

OpenStudy (anonymous):

you can write these all in exponential form it that would help? \[(6^{1/2}-3^{1/15})(6^{1/2}+3^{1/6})\]

OpenStudy (anonymous):

get a largest common denominator for the exponents and all you do is foil the number ( when you foil however, the exponents add they do not multiply